# 4D Palatini–Cartan Gravity in Hamiltonian Form

### v1

@@ 1. Introduction
In the Hamiltonian formalism, Palatini–Cartan (PC) gravity in four dimensions consists of a theory with vanishing evolution Hamiltonian and several constraints—generating internal gauge transformations and diffeomorphisms, both tangent and transversal to the Cauchy surface—see [PhysRevD.58.124029, Alexandrov_2000, BarrosESa2001, Date:2008rb, Rezende:2009sv, Bodendorfer_2013, CS2019, CCS2020] and references therein. The goal of this note is to recover this not from a study of the constraints that arise in the boundary analysis, but via the elimination of an auxiliary field that happens to be present in PC gravity (with some assumptions essentially requiring that we are working on a globally hyperbolic space–time).
@@
We focus on a space–time manifold $M$ of the form $\Sigma\times I$, where $\Sigma$ is a $3$-manifold, possibly with boundary, and $I$ is an interval. Under the assumption that the metric induced on $\Sigma\times\{t\}$ by the coframe is positive-definite $\forall t\in I$, [^1] Our result actually extends to the weaker assumption that the induced metrics are nondegenerate. we show that a part of the connection plays the role of an auxiliary field that is set to zero by its own Euler–Lagrangian (EL) equations. As a consequence, we conclude that PC theory is classically equivalent to a Hamiltonian system for fields on $\Sigma$ (the restriction of ${\boldsymbol{e}}$ and part of the restriction of ${\boldsymbol{\omega}}$) with zero Hamiltonian and several first-class constraints (the torsion-free condition and the constraints known as the momentum and Hamiltonian constraint).
@@
All this fits nicely with the study of the constraints of PC theory (see [CS2019, CCS2020] and references therein), as well as with the analogue study in Einstein–Hilbert gravity (see [MR3379262, Chapter VIII] and references therein).
In the companion paper [CCfuture], we extend this study to the BV setting and also to higher dimensions, proving that the BV PC theory is strongly equivalent (via a BV pushforward [CMR2015]) to the AKSZ [AKSZ] theory obtained in [CCS2020b] from the BFV description of the reduced phase space of PC gravity.
@@ 2.1. Local coordinates
In local coordinates, the coframe ${\boldsymbol{e}}$ has components $e_{\mu}^{a}$ with $\mu$ and $a$ ranging from $1$ to $4$. The matrix $(e_{\mu}^{a})$ is required to be invertible. The Greek index is understood as a space–time index, whereas the Latin index as an internal one. The components of the connection ${\boldsymbol{\omega}}$ are $\omega_{\mu}^{ab}$ with the condition $\omega_{\mu}^{ba}=-\omega_{\mu}^{ab}$. The curvature $F_{\boldsymbol{\omega}}$ of ${\boldsymbol{\omega}}$ has components $F_{\mu\nu}^{ab}$ and is antisymmetric both in the lower and in the upper indices. It is defined as
@@ 2.2. Intrinsic formulation
where $\Gamma$ denotes the space of sections, $\Lambda$ the exterior power, and $\Omega(M)$ the space of differential forms. We call an element of $\Omega^{k,l}(M)$ a $(k,l)$-form.
@@
The argument of the integral in (2.1) is the $(4,4)$-form obtained by taking the (nondisplayed) wedge products both in $\Lambda^{\bullet}T^{*}M$ and $\Lambda^{\bullet}\mathcal{V}$ [^3] Our convention for signs in the tensor product of these two algebras is by the total degree (Deligne convention). Namely, if $\alpha$ is a $(k,l)$-form and $\beta$ an $(r,s)$-form, we have $\alpha\beta=(-1)^{(k+l)(r+s)}\beta\alpha$. One can show (see, e.g., [CCS2020, Section 2.1]) that the space $\Omega^{4,4}(M)$ of $(4,4)$-forms is canonically identified with the space of densities on $M$, so the integrand
@@ 2.3. Relation with the Einstein–Hilbert formalism
${\mathrm{d}}_{\boldsymbol{\omega}}({\boldsymbol{e}}^{2})$ $=0,$
${\boldsymbol{e}}F_{\boldsymbol{\omega}}+\frac{\Lambda}{3!}{\boldsymbol{e}}^{3}$ $=0,$
where ${\mathrm{d}}_{\boldsymbol{\omega}}$ denotes the covariant derivative. [^5] In local coordinates, $({\mathrm{d}}_{\boldsymbol{\omega}}{\boldsymbol{e}})_{\mu\nu}^{a}=\partial_{[\mu}^{\phantom{a}}e_{\nu]}^{a}+\omega_{[\mu}^{ab}e_{\nu]}^{c}\eta_{bc}$. By the Leibniz rule, the first equation is equivalent to ${\boldsymbol{e}}{\mathrm{d}}_{\boldsymbol{\omega}}{\boldsymbol{e}}=0$, which turns out to be equivalent to ${\mathrm{d}}_{\boldsymbol{\omega}}{\boldsymbol{e}}=0$ thanks to the nondegeneracy condition on ${\boldsymbol{e}}$ [CCS2020]
@@ 3. PC gravity on a cylinder
We now specialize to the case when space–time $M$ is of the form $\Sigma\times I$, where $I=[t_{0},t_{1}]$ is an interval and $\Sigma$ is a $3$-manifold. We will interpret $I$ as a time interval, even though the results of this paper hold in more generality; in other words, we are thinking of $M$ as of a globally hyperbolic manifold. In this section, $\Sigma$ is assumed to be closed (i.e., compact without boundary).
We denote by $t$ the coordinate on $I$ and by $\partial_{\text{n}}$ the vector field $\frac{\partial}{\partial t}$. (We use the index n to stress that this is the direction “normal” to $\Sigma$.) We call a $(k,l)$-form $\underline{\boldsymbol{\alpha}}$ horizontal if it satisfies $\iota_{\partial_{\text{n}}}\underline{\boldsymbol{\alpha}}=0$, where $\iota$ denotes contraction with a vector field. We will use underlines to denote horizontal forms.
A $(k,l)$-form $\boldsymbol{\alpha}$ on $\Sigma\times I$ can be written uniquely as $\boldsymbol{\alpha}=\alpha_{\text{n}}{\mathrm{d}}t+\underline{\boldsymbol{\alpha}}$ with $\underline{\boldsymbol{\alpha}}$ horizontal. [^6] In local coordinates, $\boldsymbol{\alpha}=\alpha_{\mu_{1}\dots\mu_{k}}{\mathrm{d}}x^{\mu_{1}}\cdots{\mathrm{d}}x^{\mu_{k}}$. If we denote by $i$ the indices from $1$ to $3$ labeling coordinates on $\Sigma$ and use the index $4$ for the coordinate on $I$, we have $\underline{\boldsymbol{\alpha}}=\alpha_{i_{1}\dots i_{k}}{\mathrm{d}}x^{i_{1}}\cdots{\mathrm{d}}x^{i_{k}}$ and $\alpha_{\text{n}}=\alpha_{i_{1}\dots i_{k-1}4}{\mathrm{d}}x^{i_{1}}\cdots{\mathrm{d}}x^{i_{k-1}}$. Note that $\alpha_{\text{n}}$ is a $(k-1,l)$-form, whereas $\underline{\boldsymbol{\alpha}}$ is still a $(k,l)$-form. For the fields ${\boldsymbol{e}}$ and ${\boldsymbol{\omega}}$, we denote the decomposition by [^7] The notation ${\widehat{\underline{{\boldsymbol{\omega}}}}}$ with a hat over ${\underline{{\boldsymbol{\omega}}}}$ is for later convenience.
@@
Here $\underline{{\mathrm{d}}}_{\widehat{\underline{{\boldsymbol{\omega}}}}}$ denotes the horizontal covariant derivative with respect to the connection ${\widehat{\underline{{\boldsymbol{\omega}}}}}$, which only has horizontal components; similarly, we denote by by $\underline{{\mathrm{d}}}$ the horizontal differential. [^8] In local coordinates, this simply means we only use Latin (space) indices. For example, $(\underline{{\mathrm{d}}}_{\widehat{\underline{{\boldsymbol{\omega}}}}}\omega_{\mathrm{n}})_{i}^{ab}=\partial_{i}\omega_{\mathrm{n}}^{ab}+\omega_{i}^{ac}\omega_{\mathrm{n}}^{db}\eta_{cd}-\omega_{i}^{bc}\omega_{\mathrm{n}}^{da}\eta_{cd}$.
@@
The term $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}{\widehat{\underline{{\boldsymbol{\omega}}}}}$ is the analogue of the $p\dot{q}$ term in the Hamiltonian version of an action functional. It has however a problem: it is invariant under certain changes in ${\widehat{\underline{{\boldsymbol{\omega}}}}}$, which leads to a degeneracy (this is intimately related to the degeneracy of the boundary structure observed in [CS2019]). Namely, if we vary ${\widehat{\underline{{\boldsymbol{\omega}}}}}$ by a horizontal $(1,2)$-form $\underline{{\boldsymbol{v}}}$, we get a term $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}\underline{{\boldsymbol{v}}}$, which can be rewritten as $\frac{1}{2}\underline{{\boldsymbol{e}}}\partial_{\mathrm{n}}(\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}})-\frac{1}{2}\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}\partial_{\mathrm{n}}\underline{{\boldsymbol{e}}}$. If $\underline{{\boldsymbol{v}}}$ satisfies $\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}=0$, which actually admits nontrivial solutions, then $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}\underline{{\boldsymbol{v}}}$ vanishes. Below we show how to deal with this problem, essentially following [CCS2020] (see also [CPencyPC, Section 7] for a shorter review).
First we pick a $(0,1)$-form $\epsilon_{\text{n}}$ normalized by $(\epsilon_{\text{n}},\epsilon_{\text{n}})=-1$. We can view $\epsilon_{\text{n}}$ as a reference time-like basis vector for the internal Minkowski space $(V,\eta)$. From now on we require the field $\underline{{\boldsymbol{e}}}$ to satisfy the additional nondegeneracy condition
@@
To proceed, we now have to make the crucial assumption introduced in [CCS2020]. We denote by $g^{\Sigma}$, at each value of $t\in I$, the restriction to $\Sigma\times\{t\}$ of the metric $g$ defined by ${\boldsymbol{e}}$ (equivalently, $g^{\Sigma}$ is the tensor defined by pulling back the internal metric $\eta$ via $\underline{{\boldsymbol{e}}}$). [^9] In local coordinates, $g^{\Sigma}_{ij}=\underline{e}_{i}^{a}\underline{e}_{j}^{b}\eta_{ab}$.
@@
Under the assumption that $\underline{{\boldsymbol{e}}}$ is metric nondegenerate, it is proven in [CCS2020, Section 4.1] that there is a unique decomposition [^11] For a related analysis in the Dirac formalism, see [BarrosESa2001, Alexandrov_2000].
@@
by a uniquely determined horizontal connection ${\underline{{\boldsymbol{\omega}}}}$ and a uniquely determined horizontal $(1,2)$-form $\underline{{\boldsymbol{v}}}$ satisfying
@@
(3.6b) $\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}$ $=0,$
where $\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}$ denotes again the horizontal covariant derivative, [^12] In local coordinates, as in footnote 8, $(\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}\underline{{\boldsymbol{e}}})_{ij}^{a}=\partial_{[i}^{\phantom{a}}\underline{e}_{j]}^{a}+\underline{\omega}_{[i}^{ab}\underline{e}_{j]}^{c}\eta_{bc}$. and $\underline{\boldsymbol{\sigma}}$ is a horizontal $(1,1)$-form. Equation (3.6a) is called the structural constraint in [CCS2020]. Finally we introduce the $(0,2)$-form
@@
A simple computation, see [CCS2020], shows that (3.6b) implies $\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}\cdot\underline{{\boldsymbol{e}}}=0$. Therefore, upon integrating by parts, we may replace $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\underline{{\mathrm{d}}}_{\widehat{\underline{{\boldsymbol{\omega}}}}}\omega_{\mathrm{n}}$ with $-\omega_{\mathrm{n}}\underline{{\boldsymbol{e}}}\,\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}\underline{{\boldsymbol{e}}}$.
@@ 3.1. Getting rid of the auxiliary term
To compute $\psi_{\not=}$ more easily, we use (3.6b), which for $e_{\mu}^{a}=\delta_{\mu}^{a}$ yields, as computed in [CCS2020, Appendix A], $v_{i}^{a4}=0$, for all $i,a=1,2,3$, as well as the relations
@@
We may then choose the following $6$ independent components of $\underline{{\boldsymbol{v}}}$: $v_{1}^{12}$, $v_{1}^{13}$, $v_{1}^{23}$, $v_{2}^{12}$, $v_{2}^{13}$, $v_{3}^{12}$.
@@
Using again $v_{i}^{a4}=0$, we see that $(y_{i},y_{j})$ is equal to the Euclidean inner product $(u_{i},u_{j})$ of the $3$-vectors $u_{i}$ given by $u_{i}^{j}=v_{i}^{ij}$ with $i,j=1,2,3$ and no sum over $i$. Explicitly,

### v2

@@ 1. Introduction
In the Hamiltonian formalism, Palatini–Cartan (PC) gravity in four dimensions consists of a theory with vanishing evolution Hamiltonian and several constraints—generating internal gauge transformations and diffeomorphisms, both tangent and transversal to the Cauchy surface—see [AV98a, Ale00a, Bar01a, DKS09a, RP09a, BTT13a, CS19c, CCS21c, Mon+20a, MRC19a, Aro+00a] and references therein. The goal of this note is to recover this not from a study of the constraints that arise in the boundary analysis, but via the elimination of an auxiliary field that happens to be present in PC gravity (with some assumptions essentially requiring that we are working on a globally hyperbolic space–time).
@@
We focus on a space–time manifold $M$ of the form $\Sigma\times I$, where $\Sigma$ is a 333-manifold, possibly with boundary, and $I$ is an interval. Under the assumption that the metric induced on $\Sigma\times\{t\}$ by the coframe is positive-definite $\forall t\in I$, [^1] Our result actually extends to the weaker assumption that the induced metrics are nondegenerate. we show that a part of the connection plays the role of an auxiliary field that is set to zero by its own Euler–Lagrangian (EL) equations. As a consequence, we conclude that PC theory is classically equivalent to a Hamiltonian system for fields on $\Sigma$ (the restriction of ${\boldsymbol{e}}$ and part of the restriction of ${\boldsymbol{\omega}}$) with zero Hamiltonian and several first-class constraints (the torsion-free condition and the constraints known as the momentum and Hamiltonian constraint).
@@
All this fits nicely with the study of the constraints of PC theory (see [CS19c, CCS21c] and references therein), as well as with the analogue study in Einstein–Hilbert gravity (see [Cho15a, Chapter VIII] and references therein).
In the companion paper [CC25a], we extend this study to the BV setting and also to higher dimensions, proving that the BV PC theory is strongly equivalent (via a BV pushforward [CMR2015]) to the AKSZ [Ale+97a] theory obtained in [CCS21b] from the BFV description of the reduced phase space of PC gravity.
@@ 2.1. Local coordinates
In local coordinates, the coframe ${\boldsymbol{e}}$ has components $e_{\mu}^{a}$ with $\mu$ and $a$ ranging from 111 to 444. The matrix $(e_{\mu}^{a})$ is required to be invertible. The Greek index is understood as a space–time index, whereas the Latin index as an internal one. The components of the connection ${\boldsymbol{\omega}}$ are $\omega_{\mu}^{ab}$ with the condition $\omega_{\mu}^{ba}=-\omega_{\mu}^{ab}$. The curvature $F_{\boldsymbol{\omega}}$ of ${\boldsymbol{\omega}}$ has components $F_{\mu\nu}^{ab}$ and is antisymmetric both in the lower and in the upper indices. It is defined as
@@ 2.2. Intrinsic formulation
[…]
@@
[…]
@@ 2.3. Relation with the Einstein–Hilbert formalism
${\mathrm{d}}_{\boldsymbol{\omega}}({\boldsymbol{e}}^{2})$ =0=0,= 0 ,
${\boldsymbol{e}}F_{\boldsymbol{\omega}}+\frac{\Lambda}{3!}{\boldsymbol{e}}^{3}$ =0=0,= 0 ,
where ${\mathrm{d}}_{\boldsymbol{\omega}}$ denotes the covariant derivative. [^5] In local coordinates, $({\mathrm{d}}_{\boldsymbol{\omega}}{\boldsymbol{e}})_{\mu\nu}^{a}=\partial_{[\mu}^{\phantom{a}}e_{\nu]}^{a}+\omega_{[\mu}^{ab}e_{\nu]}^{c}\eta_{bc}$. By the Leibniz rule, the first equation is equivalent to ${\boldsymbol{e}}{\mathrm{d}}_{\boldsymbol{\omega}}{\boldsymbol{e}}=0$, which turns out to be equivalent to ${\mathrm{d}}_{\boldsymbol{\omega}}{\boldsymbol{e}}=0$ thanks to the nondegeneracy condition on ${\boldsymbol{e}}$ [CCS21c]
@@ 3. PC gravity on a cylinder
We now specialize to the case when space–time $M$ is of the form $\Sigma\times I$, where $I=[t_{0},t_{1}]$ is an interval and $\Sigma$ is a 333-manifold. We will interpret $I$ as a time interval, even though the results of this paper hold in more generality; in other words, we are thinking of $M$ as of a globally hyperbolic manifold. In this section, $\Sigma$ is assumed to be closed (i.e., compact without boundary).
[…]
@@
Here $\underline{{\mathrm{d}}}_{\widehat{\underline{{\boldsymbol{\omega}}}}}$ denotes the horizontal covariant derivative with respect to the connection $^{\widehat{\underline{{\boldsymbol{\omega}}}}}$, which only has horizontal components; similarly, we denote by by $\underline{{\mathrm{d}}}$ the horizontal differential. [^8] In local coordinates, this simply means we only use Latin (space) indices. For example, $(\underline{{\mathrm{d}}}_{\widehat{\underline{{\boldsymbol{\omega}}}}}\omega_{\mathrm{n}})_{i}^{ab}=\partial_{i}\omega_{\mathrm{n}}^{ab}+\omega_{i}^{ac}\omega_{\mathrm{n}}^{db}\eta_{cd}-\omega_{i}^{bc}\omega_{\mathrm{n}}^{da}\eta_{cd}$.
@@
The term $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}{\widehat{\underline{{\boldsymbol{\omega}}}}}$ is the analogue of the $p\dot{q}$ term in the Hamiltonian version of an action functional. It has however a problem: it is invariant under certain changes in $^{\widehat{\underline{{\boldsymbol{\omega}}}}}$, which leads to a degeneracy (this is intimately related to the degeneracy of the boundary structure observed in [CS19c]). Namely, if we vary $^{\widehat{\underline{{\boldsymbol{\omega}}}}}$ by a horizontal (1,2)(1,2)( 1 , 2 )-form $\underline{{\boldsymbol{v}}}$, we get a term $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}\underline{{\boldsymbol{v}}}$, which can be rewritten as $\frac{1}{2}\underline{{\boldsymbol{e}}}\partial_{\mathrm{n}}(\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}})-\frac{1}{2}\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}\partial_{\mathrm{n}}\underline{{\boldsymbol{e}}}$. If $\underline{{\boldsymbol{v}}}$ satisfies $\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}=0$, which actually admits nontrivial solutions, then $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}\underline{{\boldsymbol{v}}}$ vanishes. Below we show how to deal with this problem, essentially following [CCS21c] (see also [Cat23a, Section 7] for a shorter review).
First we pick a (0,1)(0,1)( 0 , 1 )-form $\epsilon_{\text{n}}$ normalized by $(\epsilon_{\text{n}},\epsilon_{\text{n}})=-1$. We can view $\epsilon_{\text{n}}$ as a reference time-like basis vector for the internal Minkowski space $(V,\eta)$. From now on we require the field $\underline{{\boldsymbol{e}}}$ to satisfy the additional nondegeneracy condition
@@
To proceed, we now have to make the crucial assumption introduced in [CCS21c]. We denote by $g^{\Sigma}$, at each value of $t\in I$, the restriction to $\Sigma\times\{t\}$ of the metric $g$ defined by ${\boldsymbol{e}}$ (equivalently, $g^{\Sigma}$ is the tensor defined by pulling back the internal metric $\eta$ via $\underline{{\boldsymbol{e}}}$). [^9] In local coordinates, $g^{\Sigma}_{ij}=\underline{e}_{i}^{a}\underline{e}_{j}^{b}\eta_{ab}$.
@@
Under the assumption that $\underline{{\boldsymbol{e}}}$ is metric nondegenerate, it is proven in [CCS21c, Section 4.1] that there is a unique decomposition [^11] For a related analysis in the Dirac formalism, see [Bar01a, Ale00a].
@@
by a uniquely determined horizontal connection ${\underline{{\boldsymbol{\omega}}}}$ and a uniquely determined horizontal (1,2)(1,2)( 1 , 2 )-form $\underline{{\boldsymbol{v}}}$ satisfying
@@
(3.6b) $\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}$ =0=0,= 0 ,
where $\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}$ denotes again the horizontal covariant derivative, [^12] In local coordinates, as in footnote 8, $(\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}\underline{{\boldsymbol{e}}})_{ij}^{a}=\partial_{[i}^{\phantom{a}}\underline{e}_{j]}^{a}+\underline{\omega}_{[i}^{ab}\underline{e}_{j]}^{c}\eta_{bc}$. and $\underline{\boldsymbol{\sigma}}$ is a horizontal (1,1)(1,1)( 1 , 1 )-form. Equation (3.6a) is called the structural constraint in [CCS21c]. Finally we introduce the (0,2)(0,2)( 0 , 2 )-form
@@
A simple computation, see [CCS21c], shows that (3.6b) implies $\underline{{\boldsymbol{e}}}\,\underline{{\boldsymbol{v}}}\cdot\underline{{\boldsymbol{e}}}=0$. Therefore, upon integrating by parts, we may replace $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\underline{{\mathrm{d}}}_{\widehat{\underline{{\boldsymbol{\omega}}}}}\omega_{\mathrm{n}}$ with $-\omega_{\mathrm{n}}\underline{{\boldsymbol{e}}}\,\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}\underline{{\boldsymbol{e}}}$.
@@ 3.1. Getting rid of the auxiliary term
To compute $\psi_{\not=}$ more easily, we use (3.6b), which for $e_{\mu}^{a}=\delta_{\mu}^{a}$ yields, as computed in [CCS21c, Appendix A], $v_{i}^{a4}=0$, for all $i,a=1,2,3$, as well as the relations
@@
We may then choose the following 666 independent components of $\underline{{\boldsymbol{v}}}$: $v_{1}^{12}$, $v_{1}^{13}$, $v_{1}^{23}$, $v_{2}^{12}$, $v_{2}^{13}$, $v_{3}^{12}$.
@@
Using again $v_{i}^{a4}=0$, we see that $(y_{i},y_{j})$ is equal to the Euclidean inner product $(u_{i},u_{j})$ of the 333-vectors $u_{i}$ given by $u_{i}^{j}=v_{i}^{ij}$ with $i,j=1,2,3$ and no sum over $i$. Explicitly,